# Weak ergodicity breaking from supersymmetry in a fermionic kinetically constrained model

### v1

@@ V Eigenstates and eigenspectrum
The Hamiltonian (1) exhibits additional structure since the $M_{1}$ Hamiltonian is not just supersymmetric but also integrable. We here first briefly review the exact solution of the $M_{1}$ model, following the presentation of Refs. [31, 32]. The Bethe ansatz parametrizes the eigenvalues and eigenstates in terms of parameters $\mu_{j}$ ($j=1,2,\dots,f$). The eigenstate $|\psi\rangle=|\mu_{1},\mu_{2},\dots,\mu_{f}\rangle$ corresponding to a set of parameters is expressed in terms of excitations created on top of the vacuum state $|0\rangle$ as
@@
One can view $g$ as the bare scattering matrix describing the phase shift when two excitations are interchanged. Since $g(\mu_{j},\mu_{j})=-1$, all coefficients are required to be distinct, since otherwise the wave function vanishes. An exception holds for $\mu_{j}=\exp(\pm i\pi/3)$, for which both the numerator and denominator vanish. We can define $g(e^{i\pi/3},e^{i\pi/3})=g(e^{-i\pi/3},e^{-i\pi/3})=1$ without affecting the solvability [43]. The phase shift (23) more generally simplifies for $\mu_{j}=\exp(\pm i\pi/3)$, since
@@
With these definitions, the state (20) is an eigenstate of the $M_{1}$ Hamiltonian with eigenvalue
@@
The momentum $p$ of an eigenstate is encoded through the relation $e^{ip}=\mu_{1}\mu_{2}\dots\mu_{f}$. The Bethe equations directly reflect the underlying supersymmetry of the model. If $\{\mu_{1},\mu_{2},\dots,\mu_{f}\}$ is a solution with all coefficients different from unity, then $\{\mu_{1},\mu_{2},\dots,\mu_{f},1\}$ is a solution with the same energy and momentum, which leads to the supersymmetry-induced degeneracies between eigenstates with $f$ and $f+1$ fermions.‘
@@
Additionally, we note that if $\{\mu_{1},\mu_{2},\dots,\mu_{f}\}$ parametrizes a solution with momentum $p$, then $\{\mu_{1}^{-1},\mu_{2}^{-1},\dots,\mu_{f}^{-1}\}$ parametrizes a solution with momentum $-p$. Within the momentum zero and $\pi$ sectors, one can single out solutions for which $\{\mu_{1},\mu_{2},\dots,\mu_{f}\}$ is not equal to $\{\mu_{1}^{-1},\mu_{2}^{-1},\dots,\mu_{f}^{-1}\}$. These solutions are not invariant under spatial inversion, but inversion-symmetric and inversion-antisymmetric eigenstates can be constructed as $|\mu_{1},\mu_{2},\dots,\mu_{f}\rangle\pm|\mu_{1}^{-1},\mu_{2}^{-1},\dots,\mu_{f}^{-1}\rangle$ by taking the plus or minus sign, respectively. The counting of the number of inversion-symmetric and inversion-antisymmetric eigenstates in constrained models has been considered recently in Ref.  [44].
@@
Heisenberg model at root of unity. The integrability of this model directly relates to the integrability of the Heisenberg XXZ model since the $M_{1}$ Hamiltonian can be mapped to the Heisenberg XXZ Hamiltonian under a system-size changing transformation [45, 31]. This transformation consists of replacing an occupied site together with the left-neighboring empty site by an occupied site, and the remaining empty sites by empty sites. Similar system size changing transformations were also recently discussed in Ref. [46]. The resulting Heisenberg Hamiltonian has anisotropy $\Delta=1/2$ and twisted boundary conditions. This anisotropy corresponds to the Heisenberg model with anistropy at ‘roots of unity’, for which the Bethe ansatz is notoriously subtle (see, e.g. Refs. [47, 48] for a detailed discussion). We mainly note that this model exhibits an exponential amount of degeneracies, in addition to the degeneracies originating from supersymmetry, since it supports a representation of the loop algebra $\mathfrak{sl}_{2}$ [47, 48, 49, 50, 51].
@@
Sub-thermal entangled eigenstates. The PXP model is known to have exact matrix product eigenstates with system size-independent bond dimensions at energies zero and ±$\pm\sqrt{2}$ [40]. When imposing the requirement that two occupied sites are separated by at least $n$ empty sites, similar exact matrix product eigenstates at energies zero and ±$\pm\sqrt{q}$ for a number of integers $q\leq n+1$ have been identified [41]. Because of the insensitivity of the bond dimension to the system size, these special eigenstates are area-law entangled, such that they can be thought of as quantum many-body scars. Here we show that, in analogy, the PXP-like fermion model hosts eigenstates with energy ±$\pm\sqrt{q}$ for some integers $q$, which obey sub-volume law entanglement scaling. These states correspond to Bethe states with integer eigenvalues for the $M_{1}$ model. Various states with integer eigenvalue appear throughout the spectrum of the $M_{1}$ model, and in Table 1 we list these eigenvalues for system size $N=12$ for each fermion number $f$. While such integer eigenvalues have been previously observed in the $M_{1}$ model [52], we are not aware of any discussion of their origin or the corresponding solutions of the Bethe equations.
@@
While Bethe states can generally be written as an MPS, the bond dimension typically grows with system size, leading to volume-law entangled eigenstates [53]. We here identify special classes of Bethe states from the solutions to the Bethe equations (26), corresponding to integer eigenvalues, and show how these states can be recast as an area-law entangled MPS. Whenever a parameter $\mu_{j}$ equals $\mu_{j}=\exp(\pm i\pi/3)$, this parameter will contribute an integer to the energy (25) since $\mu_{j}+1/\mu_{j}=1$.
@@
We here first focus on special states in which all parameters are identical and satisfy $\mu_{j}=\exp(\pm i\pi/3)$. The eigenvalues (25) can be directly checked to satisfy $E=N-f$. The Bethe equations simplify to a single equation,
@@
For system sizes that are an even (odd) integer multiple of 3, this equation is satisfied whenever $f$ is odd (even). The corresponding eigenstates (20) simplify since
@@
As apparent from Table 1, the model supports additional eigenstates with integer eigenvalues beyond these states with $E=N-f$ (and their superpartners, which have the same energy but a fermion number differing by one). A large number of these states can be understood as ‘simple’ solutions to the Bethe equations that are dressed by additional parameters $\exp(\pm i\pi/3)$. Consider e.g. $f=1$, for which the dynamics was already argued to reduce to that of a free particle, and the Bethe equation for the single parameter $\mu$ reduces to
@@
Consider now a solution $(\mu_{1},\mu_{2},\dots\mu_{f})$ to the Bethe equations (26) for fixed $N$ and momentum $p$. Then
$(\mu_{1},\mu_{2},\dots\mu_{f},\underbrace{e^{i\pi/3},\dots,e^{i\pi/3}}_{n_{+}}\,,\underbrace{e^{i\pi/3},\dots,e^{i\pi/3}}_{n_{-}})$ (33)
@@ VI Conclusions and outlook
In summary, we studied weak ergodicity breaking in the context of $\mathcal{N}=2$ supersymmetric quantum mechanics. We have shown how supersymmetry naturally induces non-ergodic dynamics by introducing a fermionic equivalent of the PXP model, which is closely related to the supersymmetric $M_{1}$ model on a chain. For this model, we have established that the quench dynamics for properly chosen initial states display (exact or approximate) periodic revivals, in analogy to what can be observed for the PXP model and more generally for weak ergodicity breaking in the context of quantum many-body scars. The necessary time scales for these revivals can be analytically obtained, and we find that these scale as the square root of system size. We identified special many-body eigenstates, with eigenvalues at (plus or minus) square roots of integers, that allowing for a matrix product state description with small bond dimension and hence exhibiting sub-thermal entanglement scaling. We argued that these special eigenstates can be thought of as quantum many-body scars.
Our work provides a number of suggestions for further investigations. The $M_{1}$ model can be generalized and deformed in various ways without breaking the supersymmetry and/or the integrability. For example, when replacing the constraint forbidding two neighboring sites to be occupied simultaneously by a constraint forbidding sites up to a distance $k$ apart from an occupied site, an $M_{k}$ model results [32]. The $M_{2}$ model is, like the $M_{1}$ model, known to be supersymmetric and integrable [56]. An arguably interesting open question is how our results translate to such generalizations and deformations of the $M_{1}$ model. Next, it is natural to ask if interpolations between the PXP and PXP-like fermion models provide new insights on the origin of quantum many-body scars. Such interpolations can be obtained, for example, by replacing the fermions in the PXP-like fermion model by proper anyons. On the level of integrability, this work opens up ways of using the toolbox of integrability to study dynamics in the class of introduced models. Finally, while the eigenstates of this PXP-like fermion model can be exactly expressed as (linear combinations of two) Bethe-ansatz states, it is unclear if and how the conserved charges underlying integrability can be constructed for arbitrary values of the chemical potential.

### v2

@@ V Eigenstates and eigenspectrum
The Hamiltonian $H_{\text{PXP}}$ of Eq. (1) exhibits additional structure since the $M_{1}$ Hamiltonian is not just supersymmetric but also integrable. We here first briefly review the exact solution of the $M_{1}$ model, following the presentation of Refs. [32, 33]. The Bethe ansatz parametrizes the eigenvalues and eigenstates in terms of parameters $\mu_{j}$ ($j=1,2,\dots,f$). The eigenstate $|\psi\rangle=|\mu_{1},\mu_{2},\dots,\mu_{f}\rangle$ corresponding to a set of parameters is expressed in terms of excitations created on top of the vacuum state $|0\rangle$ as
@@
One can view $g$ as the bare scattering matrix describing the phase shift when two excitations are interchanged. Since $g(\mu_{j},\mu_{j})=-1$, all coefficients are required to be distinct, since otherwise the wave function vanishes. An exception holds for $\mu_{j}=\exp(\pm i\pi/3)$, for which both the numerator and denominator vanish. We can define $g(e^{i\pi/3},e^{i\pi/3})=g(e^{-i\pi/3},e^{-i\pi/3})=1$ without affecting the solvability [45]. The phase shift of Eq. (22) more generally simplifies for $\mu_{j}=\exp(\pm i\pi/3)$, since
@@
With these definitions, the state $|\psi\rangle$ as introduced in Eq. (19) is an eigenstate of the $M_{1}$ Hamiltonian with eigenvalue
@@
The momentum $p$ of an eigenstate is encoded through the relation $e^{ip}=\mu_{1}\mu_{2}\dots\mu_{f}$. The Bethe equations directly reflect the underlying supersymmetry of the model. If $\{\mu_{1},\mu_{2},\dots,\mu_{f}\}$ is a solution with all coefficients different from unity, then $\{\mu_{1},\mu_{2},\dots,\mu_{f},1\}$ is a solution with the same energy and momentum, which leads to the supersymmetry-induced degeneracies between eigenstates with $f$ and $f+1$ fermions.
@@
Additionally, we note that if $\{\mu_{1},\mu_{2},\dots,\mu_{f}\}$ parametrizes a solution with momentum $p$, then $\{\mu_{1}^{-1},\mu_{2}^{-1},\dots,\mu_{f}^{-1}\}$ parametrizes a solution with momentum p-p. Within the momentum zero and $\pi$ sectors, one can single out solutions for which $\{\mu_{1},\mu_{2},\dots,\mu_{f}\}$ is not equal to $\{\mu_{1}^{-1},\mu_{2}^{-1},\dots,\mu_{f}^{-1}\}$. These solutions are not invariant under spatial inversion, but inversion-symmetric and inversion-antisymmetric eigenstates can be constructed as $|\mu_{1},\mu_{2},\dots,\mu_{f}\rangle\pm|\mu_{1}^{-1},\mu_{2}^{-1},\dots,\mu_{f}^{-1}\rangle$ by taking the plus or minus sign, respectively. The counting of the number of inversion-symmetric and inversion-antisymmetric eigenstates in constrained models has been considered recently in Ref.  [46].
V.1 Heisenberg model at root of unity
@@
The integrability of this model directly relates to the integrability of the Heisenberg XXZ model since the $M_{1}$ Hamiltonian can be mapped to the Heisenberg XXZ Hamiltonian under a system-size changing transformation [47, 32]. This transformation consists of replacing an occupied site together with the left-neighboring empty site by an occupied site, and the remaining empty sites by empty sites. Similar system size changing transformations were also recently discussed in Ref. [48]. The resulting Heisenberg Hamiltonian has anisotropy $\Delta=1/2$ and twisted boundary conditions. This anisotropy corresponds to the Heisenberg model with anistropy at ‘roots of unity’, for which the Bethe ansatz is notoriously subtle (see, e.g. Refs. [49, 50] for a detailed discussion). We mainly note that this model exhibits an exponential amount of degeneracies, in addition to the degeneracies originating from supersymmetry, since it supports a representation of the loop algebra $\mathfrak{sl}_{2}$ [49, 50, 51, 52, 53].
V.2 Sub-thermal entangled eigenstates
@@
The PXP model is known to have exact matrix product eigenstates with system size-independent bond dimensions at energies zero and $\pm\sqrt{2}$ [42]. When imposing the requirement that two occupied sites are separated by at least $n$ empty sites, similar exact matrix product eigenstates at energies zero and $\pm\sqrt{q}$ for a number of integers $q\leq n+1$ have been identified [43]. Because of the insensitivity of the bond dimension to the system size, these special eigenstates are area-law entangled. Here we show that, in analogy, the PXP-like fermion model hosts eigenstates with energy $\pm\sqrt{q}$ for some integers $q$, which obey sub-volume law entanglement scaling. These states correspond to Bethe states with integer eigenvalues for the $M_{1}$ model. Various states with integer eigenvalue appear throughout the spectrum of the $M_{1}$ model, and in Table 1 we list these eigenvalues for system size $N=12$ for each fermion number $f$. While such integer eigenvalues have been previously observed in the $M_{1}$ model [54], we are not aware of any discussion of their origin or the corresponding solutions of the Bethe equations.
@@
While Bethe states can generally be written as an MPS, the bond dimension typically grows with system size, leading to volume-law entangled eigenstates [55]. We here identify special classes of Bethe states from the solutions to the Bethe equations as given in Eq. (25), corresponding to integer eigenvalues, and show how these states can be recast as an area-law entangled MPS. Whenever a parameter $\mu_{j}$ equals $\mu_{j}=\exp(\pm i\pi/3)$, this parameter will contribute an integer to the energy $E$ of Eq. (24) since $\mu_{j}+1/\mu_{j}=1$.
@@
We here first focus on special states in which all parameters are identical and satisfy $\mu_{j}=\exp(\pm i\pi/3)$. The eigenvalues given in Eq. (24) can be directly checked to satisfy E=NfE=N-f. The Bethe equations simplify to a single equation,
@@
For system sizes that are an even (odd) integer multiple of 3, this equation is satisfied whenever $f$ is odd (even). The corresponding eigenstates of Eq. (19) simplify since
@@
As apparent from Table 1, the model supports additional eigenstates with integer eigenvalues beyond these states with E=NfE=N-f (and their superpartners, which have the same energy but a fermion number differing by one). A large number of these states can be understood as ‘simple’ solutions to the Bethe equations that are dressed by additional parameters $\exp(\pm i\pi/3)$. Consider e.g. $f=1$, for which the dynamics was already argued to reduce to that of a free particle, and the Bethe equation for the single parameter $\mu$ reduces to
@@
Consider now a solution $(\mu_{1},\mu_{2},\dots\mu_{f})$ to the Bethe equations given in Eq. (25) for fixed $N$ and momentum $p$. Then
$(\mu_{1},\mu_{2},\dots\mu_{f},\underbrace{e^{i\pi/3},\dots,e^{i\pi/3}}_{n_{+}}\,,\underbrace{e^{-i\pi/3},\dots,e^{-i\pi/3}}_{n_{-}})$ (32)
@@ VI Conclusions and outlook
In summary, we studied the dynamics of kinetically constrained models constructed from a supersymmetric algebra. For Hamiltonians of the form of Eq. (4) the dynamics was argued to be generally non-ergodic, and periodic revivals can be anticipated based on the underlying doublet structure induced by supersymmetry. We illustrated this construction by introducing a fermionic equivalent of the PXP model with an adjustable chemical potential, closely related to the supersymmetric $M_{1}$ model on a chain. For this model, we established that the quench dynamics for properly chosen initial states display (exact or approximate) periodic revivals, in analogy to what can be observed for the PXP model. The necessary time scales for these revivals can be analytically obtained, and we find that these scale as the square root of system size. Utilizing the Bethe-integrability of the $M_{1}$ model, we next identified special many-body eigenstates, with eigenvalues at (plus or minus) square roots of integers, that allowing for a matrix product state description with small bond dimension and hence exhibiting sub-thermal entanglement scaling.
Our work provides a number of suggestions for further investigations. It is natural to ask if interpolations between the PXP and PXP-like fermion models provide new insights on the origin of quantum many-body scars in the former. Such interpolations can be obtained, for example, by replacing the fermions in the PXP-like fermion model by proper anyons. The $M_{1}$ model can also be generalized and deformed in various ways without breaking the supersymmetry and/or the integrability. For example, when replacing the constraint forbidding two neighboring sites to be occupied simultaneously by a constraint forbidding sites up to a distance $k$ apart from an occupied site, an $M_{k}$ model results [33]. The $M_{2}$ model is, like the $M_{1}$ model, known to be supersymmetric and integrable, while $M_{k}$ models are supersymmetric for all $k$ [58]. The $M_{2}$ model has recently gained attention since it additionally corresponds to “free fermions in disguise” [59]. An interesting open question is how our results translate to such generalizations and deformations of the $M_{1}$ model, and to supersymmetric fermion lattice models in general. Shifting the focus to integrability, this work opens up ways of using the toolbox of integrability to study dynamics in the class of introduced models. Finally, while the eigenstates of this PXP-like fermion model can be exactly expressed as (linear combinations of two) Bethe-ansatz states, it is unclear if and how the conserved charges underlying integrability can be constructed for arbitrary values of the chemical potential.